Ammonium nitrate, a common fertilizer, is used as an explosive in fireworks and by terrorists. It was the material used in the tragic explosion at the Oklahoma City federal building in 1995. How many liters of gas at 307°C and 1.00 atm are formed by the explosive decomposition of 80.3 kg of ammonium nitrate to nitrogen, oxygen, and water vapor? Do not answer in scientific notation.

1
Expert's answer
2021-03-08T06:04:20-0500

2NH4NO3 = 2N2 + O2 + 4H2O

80.3kg = 80300g

2 mole ammonium nitrate = 2 moles nitrogen gas

2 × 80g NH4NO3 = 2mole N2

1g NH4NO3 = 2/160 mole nitrogen gas

80300g 2NH4NO3 = 2 × 80300/160 mole N2

= 1003.75mol


2 mole NH4NO3 = 1 mole oxygen gas

2×80g NH4NO3 = 1 mole oxygen gas

1g NH4NO3 = 1/160 mole oxygen gas

80300g NH4NO3 = 80300/160 mole oxygen gas

= 501.875 mole oxygen gas


2 mole NH4NO3 = 4 mole H20

2×80 g NH4NO3 = 4mole H20

1g NH4NO3 = 4/160 mole H2O

80300g NH4NO3 = 4×80300/160 mole H2O

= 2007.5 mole of H20


PV=nRT . T= 307°C = 34K

V = nRT / P


For nitrogen V= 1003.75 × 0.082 × 34 / 1 = 2798.46L

For oxygen V = 501.875 × 0.082 × 34 = 1399.23L

For water vapour V = 2007.5 × 0.082 × 34 = 5596.91L


Total volume = 2798.46L + 1399.23L + 5596.91L

= 9794.6L